# ExPLoRA: Parameter-Efficient Extended Pre-Training to Adapt Vision Transformers under Domain Shifts

### v2

@@ 5 Method
To learn $\Delta_{D_{T}}$, we propose ExPLoRA (i.e. Extended Pre-training with LoRA), a method that efficiently adapts pre-trained ViTs for new target domains $D_{T}$, described in algorithm 1.
In terms of notation, D-[L]-$r64$ refers to a ViT initialized with DinoV2 weights (denoted by D), where ${\mathcal{U}}=\{\texttt{L}\}$, and LoRA rank 64 is applied to the $Q,V$ matrices of attention layers in ${\mathcal{L}}\setminus{\mathcal{U}}$. Thus, $\Delta_{D_{T}}$ comprises of all weights in ${\mathcal{U}}$, LoRA matrices in ${\mathcal{L}}\setminus{\mathcal{U}}$, and normalization layers. For ${\mathcal{U}}=\{\texttt{L}\}$, $\Delta_{D_{T}}$ consists of only 5% of the original ViT parameters. As we show in section 6, our extended pre-training approach can match or even outperform full pre-training on new domains from scratch.
ExPLoRA for DinoV2
We initialize a ViT-L with $W_{D_{S}}$ from the DinoV2 ViT-L encoder, without registers (Darcet et al., 2023). Since the DinoV2 pre-trained checkpoints don’t contain the Dino or iBOT linear heads, we initialize a shared Dino-iBOT linear head from scratch. This shared head is fully trained during extended pre-training, adding only a minimal number of trainable parameters.
ExPLoRA for MAE
We initialize a ViT-L with $W_{D_{S}}$ from the MAE ViT-L encoder. Since MAE provides the pre-trained decoder, we use these weights to initialize our MAE decoder He et al. (2022). During extended pre-training, in addition to the ExPLoRA recipe in algorithm 1, we apply LoRA with rank $r^{\prime}$ on the $Q,V$ matrices of each attention layer in the frozen decoder. Note that the LoRA rank $r^{\prime}$ may differ from the LoRA rank $r$ used in the ViT encoder (section A.4). All other decoder weights, apart from the layer-normalization layers, are kept frozen. No block is fully unfrozen in the MAE decoder, as it will be discarded after extended pre-training. This helps to minimize the number of additional parameters trained in the decoder.
ExPLoRA for Multi-Spectral Inputs
For the multi-spectral ViT introduced by SatMAE we need to additionally unfreeze the positional encoding and the patch embedding weights for each group of channels. These cannot be initialized from $W_{D_{S}}$, as $W_{D_{S}}$ is trained on RGB inputs, whereas multi-spectral inputs can have more or different channels. As part of $\Delta_{D_{T}}$ in algorithm 1, the positional encodings and patch embeddings for multi-spectral data are adapted during extended pre-training. Aside from this, the approach remains unchanged from that of DinoV2 or MAE described earlier.
Storage Considerations
After running ExPLoRA, we receive a new unsupervised model $W_{D_{T}}^{\ast}=W_{D_{S}}+\Delta_{D_{T}}$ for the target domains $D_{T}$. Any components that are not part of the ViT encoder (eg: the Dino linear head or the MAE decoder) are discarded. Post-ExPLoRA, only $\Delta_{T}$, consisting of 1-2 unfrozen ViT blocks, LoRA matrices, and layer-normalization weights, are stored for each $D_{T}$. Like LoRA and other PEFT methods, ExPLoRA significantly reduces additional storage requirements compared to fully training $W_{D_{T}}$ from scratch.
@@
After extended pre-training with ExPLoRA, the output weights $W_{D_{T}}^{\ast}$ behave as any fully pre-trained model $W_{D_{T}}$. We can now use the ViT for feature extraction, PEFT, or fine-tuning as desired. For instance, we could initialize a linear head for classification or a decoder for segmentation, either of which is fully trainable. We can then freeze all ViT weights and apply LoRA-$r8$ on the $Q,V$ matrices of the attention layers (or use another PEFT method). Lastly, we use supervised fine-tuning on each labeled dataset ${\mathcal{Y}}_{d_{T}}$ to train the unfrozen parameters $\Delta_{d_{T}}$. This yields our final model $W_{d_{T}}$ (eq. 4), which can be used for classification, segmentation, detection etc.
@@ 6 Experiments
Our experimental results consist of a case study on satellite imagery (section 6.1), with an ablation study in section 6.1.2 and analysis in section 6.3. We evaluate on multiple downstream tasks in sections 6.1.3, 6.1.4 and 6.2. Additional experiments and ablations are provided in appendix A and training hyperparameter and compute configurations are mentioned in appendix B. Our results achieve a new SoTA top 1 accuracy of 79.2% ($\uparrow$1.4%) on the competitive fMoW-RGB benchmark, outperforming fully pre-trained and fine-tuned models while using 6% of the ViT encoder parameters. We also achieve a $\uparrow$8.2improvement in linear probing accuracy on the same dataset. Across other satellite datasets, we match fully-pretrained prior state-of-the-art methods, and demonstrate competitive performance on WiLDS benchmark datasets as well.
@@ 6.1 Case Study: Satellite Imagery
We examine satellite images given their importance towards multiple societal applications (Broader Impact) and since they represent a significant domain shift from natural images. There is a large and growing body of research on developing foundational models for satellite imagery from scratch Ayush et al. (2021); Cong et al. (2022); Reed et al. (2023); Tang et al. (2024), thus presenting a good benchmark for ExPLoRA.
@@ 6.1.1 RGB Satellite Images
We first consider the functional map of the world (fMoW) dataset of high-resolution satellite images, each paired with one of 62 classification labels (Christie et al., 2018). fMoW is used as a benchmark for many satellite-image foundation models (Cong et al., 2022; Reed et al., 2023).
Model Arch. PEFT
Pre-train
@@
Fine-tune
#Params
Top 1 Acc.
ScaleMAE [51] ViT-L Full 303.3M 303.3M 77.80
SatMAE [15] ViT-L Full 303.3M 303.3M 77.78
SatMAE [15] ViT-L LoRA-$r8$ [28] 303.3M 0.8M 76.10
MAE [26] ViT-L LoRA-$r8$ [28] - 0.8M 76.21
MAE [26] ViT-L DVPT-10 [30] - 0.4M 72.35
MAE [26] ViT-L GVPT-100 [60] - 0.4M 70.86
MAE [26] ViT-L SA2VP [47] - 1.1M 73.55
MAE [26] ViT-L AdaLoRA-$r8$ [62] - 1.2M 75.25
M-[L]-$r64$ ViT-L LoRA-$r8$ [28] 18.7M 0.8M 76.55
DinoV2 [45] ViT-L LoRA-$r8$ [28] - 0.8M 78.08
DinoV2 [45] ViT-L BOFT-$b2m8$ [40] - 0.9M 72.40
DinoV2 [45] ViT-L VPT-100 [30] - 0.4M 77.29
DinoV2 [45] ViT-L GVPT-100 [60] - 0.4M 76.22
DinoV2 [45] ViT-L AdaLoRA-$r8$ [62] - 1.2M 78.87
DinoV2 [45] ViT-L SA2VP [47] - 1.1M 77.53
D-[L]-$r64$ ViT-L SA2VP [47] 18.7M 1.1M 78.51
D-[L]-$r64$ ViT-L LoRA-$r8$ [28] 18.7M 0.8M 79.15
Table 1: Results on the fMoW-RGB validation dataset. The “Pre-train #Params" and “Fine-tune #Params" refer to the trainable parameters of the ViT encoder required on the new domain, i.e. satellite images. M-[L]-$r64$ and D-[L]-$r64$ refer to ExPLoRA models initialized with MAE and DinoV2 weights, respectively (section 5).
We compare our results in table 1 against both prior fully pre-trained SoTA foundation models as well as PEFT techniques applied on ViTs pre-trained with MAE and/or DinoV2 weights. Our results demonstrate that D-ExPLoRA-$\tt{[L]}$-$r64$ is SoTA in terms of fMoW-RGB average accuracy at 79.15%. ExPLoRA outperforms techniques that require fully pre-training ViTs on fMoW while using 6% of the original ViT encoder parameters.
ExPLoRA-initializations with LoRA fine-tuning outperform other unsupervised initializations paired with PEFT techniques by 1-3%, including SoTA matrix-adaptation methods like AdaLoRA (Zhang et al., 2023b), BOFT (Liu et al., 2023), and VPT-based approaches such as GVPT (Yoo et al., 2023) and SA2VP (Pei et al., 2024). Additionally, applying SA2VP to ExPLoRA-initialized ViTs improves performance over a DinoV2 initialization by  1%, showcasing ExPLoRA’s compatibility with other PEFT methods and its versatility as an initialization for new domains.
6.1.2 Ablation study
Method Arch.
@@ 6.1.2 Ablation study
Table 2: Linear-probing on fMoW-RGB. The first four rows fully pre-train on the dataset. $\dagger$ denotes concatenating features from the last 4 ViT blocks. All other rows use the features of the last ViT block.
@@
Next, we investigate whether high LoRA ranks used on all ViT layers (i.e. all attention and MLP matrices, not just $Q,V$) is beneficial. Surprisingly, this significantly harms learning (row 4). In fact, it is much less effective than using just LoRA-$r256$ on the $Q,V$ matrices of all ${\mathcal{L}}$ blocks (row 3). However, both rows 3 and 4 are much less parameter-efficient than ExPLoRA (rows 5-7, 10-12).
@@
Lastly, we investigate the impact of LoRA rank on ExPLoRA. Changing the rank from 8 to 32 has a small improvement ($\uparrow 0.09\%$), but changing from 32 to 64 brings about a much larger improvement ($\uparrow 1.08\%$), with only a relatively small increase in trainable parameters. This demonstrates that higher ranks are necessary during pre-training for effective learning on the new domain. Further ablations on data efficiency (A.3), MAE decoder rank (A.4), and ViT backbone size (A.5) are in appendix A.
@@ 6.1.3 Multi-Spectral Satellite Images
Dataset
Next, we consider the fMoW-Sentinel dataset, a large dataset of Sentinel-2 images used in Cong et al. (2022). Each image consists of 13 spectral bands and is paired with one of 62 classes.
With fMoW-Sentinel, we evaluate the challenge of transferring from natural images to multi-spectral, low-resolution satellite images– a much harder task than fMoW-RGB due to the absence of sensor data beyond RGB bands in $D_{S}$. We use the SatMAE group-channel ViT-L model from Cong et al. (2022), initialized with MAE weights from natural images. Since the patch embedding layers differ, we unfreeze and train them from scratch during algorithm 1, adding minimal parameter overhead.
Model Backbone PEFT
Pre-train
@@
Fine-tune
#Params
Top 1 Acc.
ImgNet-Supervised ResNet152 Full 60.3M 60.3M 54.46
MAE [26] ViT-L Full - 303.3M 51.61
SatMAE [15] ViT-L Full 303.3M 303.3M 61.48
MAE [26] ViT-L LoRA-r8 - 0.8M 46.97
SatMAE [15] ViT-L LoRA-r8 303.3M 0.8M 59.48
MAE-[1,2,L-1,L] ViT-L LoRA-r8 51.5M 0.8M 54.12
M-ExPLoRA-[L]-$r32$ ViT-L LoRA-r8 16.2M 0.8M 51.84
M-ExPLoRA-[1,L]-$r32$ ViT-L LoRA-r8 29.7M 0.8M 60.15
Table 4: Results on the fMoW-Sentinel validation set. The “Pre-train #Params" and “Fine-tune #Params" refer to the trainable parameters required on the new domain, i.e. multi-spectral satellite images. “MAE-[1,2,L-1,L]" refers to initializing the group-channel SatMAE model with MAE weights, unfreezing blocks 1,2,23,24 for ViT-L, and then continuing pre-training on fMoW-Sentinel.
Table 4 highlights the difficulty of this domain transfer. Fully fine-tuning from MAE weights leads to a nearly 10% drop in accuracy (row 2), and LoRA tuning from MAE alone performs even worse (row 4). Simply unfreezing four transformer blocks during pre-training (row 6) fails to close the gap. However, ExPLoRA, with ${\mathcal{U}}=\{\tt{1,L}\}$, yields excellent results, even outperforming full pre-training from scratch for LoRA fine-tuning (row 5 vs. last row). This demonstrates ExPLoRA’s effectiveness in bridging large domain gaps while using just a fraction (here, <10%) of the parameters.
6.1.4 Additional Satellite Datasets
We perform extensive experiments on downstream satellite datasets, with further results in A.1.
fMoW-Temporal
Each input is a sequence of up to 3 fMoW-RGB (Christie et al., 2018) images of the same location, distributed temporally, and paired with one of 62 classes. Since the inputs are now temporal sequences, we initialize the temporal MAE architecture from Cong et al. (2022) with MAE weights, and pre-train on ${\mathcal{X}}_{D_{T}}$ with ${\mathcal{U}}=\tt[L]$ and LoRA rank 32. ExPLoRA then outperforms temporal SatMAE for PEFT (table 6), demonstrating successful transfer learning at a fraction of the pre-training parameters.
SpaceNet-v1
RESISC-45
The RESISC-45 (Cheng et al., 2017) benchmark dataset consists of 31,500 satellite images of varying resolution (0.2m-30m GSD), with 45 classes. The data is split into 25,200 training and 6,300 validation images, as per Reed et al. (2023). In table 6, our D-ExPLoRA pre-trained on only high-resolution fMoW-RGB images achieves SoTA results of 97.32% on multi-resolution RESISC-45 images, with just linear-probing. Since we use the same pre-trained model as in the last row of table 1, we demonstrate successful transfer learning from ExPLoRA pre-training, without requiring any additional modifications for scale-aware representation learning (Reed et al., 2023).
@@ 6.1.4 Additional Satellite Datasets
Table 5: fMoW-Temporal validation set results
@@
Table 6: SpaceNet and Resisc-45 validation set results
6.2 WiLDS Datasets
We test ExPLoRA on the WILDS (Koh et al., 2021) benchmark, specifically on Camelyon17 (Bandi et al., 2018), iWildcam (Beery et al., 2020) and GlobalWheat David et al. (2020; 2021) datasets, representing domain transfers to medical, wildlife, and agricultural imagery, respectively.
@@ 6.2 WiLDS Datasets
Table 9: Object detection results on the validation set of GlobalWheat. AP and AR stand for average precision and average recall, respectively.

### v3

@@ 5 Method
To learn $\Delta_{T}$, we propose ExPLoRA (i.e. Extended Pre-training with LoRA), a method that efficiently adapts a pre-trained ViT to a new target domain $T$, in algorithm 1.
Concretely, we learn $\Delta_{T}$ as follows:
$\Delta_{T}=\operatorname*{arg\,min}_{\theta\in\Theta({\mathcal{U}},r)}\left(\min_{\psi}\sum_{{\mathbf{x}}\in{\mathcal{X}}_{T}}{\mathcal{C}}_{S}\left(g_{\psi}\left(f_{\theta}({\mathbf{x}};W_{S})\right),{\mathbf{x}}\right)\right)$ (5)
where $f_{\theta}(\cdot;W_{S})$ is a ViT feature encoder parameterized by $\theta$ with initialization $W_{S}$, $g_{\psi}$ is a learnable decoder parameterized by $\psi$ (e.g., MAE/DinoV2 decoder), ${\mathcal{C}}_{S}$ is the unsupervised loss that was used for $W_{S}$ (e.g., reconstruction loss), and $\Theta({\mathcal{U}},r)$ restricts the parameter space to full-rank updates for blocks in ${\mathcal{U}}$ and LoRA rank-$r$ updates in ${\mathcal{L}}\setminus{\mathcal{U}}$. As we show in section 6, ExPLoRA can even outperform full pre-training on new domains from scratch.
In terms of notation, D-[L]-$r64$ refers to a ViT initialized with DinoV2 weights, where ${\mathcal{U}}=\{\texttt{L}\}$, and LoRA rank 64 is applied to the $Q,V$ matrices of attention layers in ${\mathcal{L}}\setminus{\mathcal{U}}$. Thus, $\Delta_{T}$ comprises of all weights in ${\mathcal{U}}$, LoRA matrices in ${\mathcal{L}}\setminus{\mathcal{U}}$, and layer normalization parameters of all blocks.
For DinoV2, we initialize a ViT-L with $W_{S}$ from DinoV2’s ViT encoder, without registers (Darcet et al., 2023). Since DinoV2 pre-trained checkpoints don’t include Dino or iBOT linear heads, we initialize a shared head $g_{\psi}$ from scratch, which is fully trained during extended pre-training.
For MAE, we initialize a ViT-L with $W_{S}$ from the MAE ViT-L encoder and use provided weights to initialize decoder $g_{\psi}$ (He et al., 2022). During extended pre-training, beyond the ExPLoRA recipe in algorithm 1, we apply LoRA with rank $r^{\prime}$ on the $Q,V$ matrices of each attention layer in $g_{\psi}$. The LoRA rank $r^{\prime}$ may differ from rank $r$ used in the ViT encoder $f_{\theta}$ (section B.4). All other decoder weights except layer-normalization are frozen, with no block fully unfrozen in $g_{\psi}$ to minimize trainable parameters.
For the multi-spectral ViT introduced by SatMAE we need to additionally unfreeze the positional encoding and patch embedding weights for each group of channels, as part of $\Theta({\mathcal{U}},r)$ in eq. 5. These cannot be initialized from $W_{S}$, as $W_{S}$ is trained on RGB inputs, whereas multi-spectral inputs can have more or different channels.
@@
After running ExPLoRA, we receive a new unsupervised model $W_{T}^{\ast}=W_{S}+\Delta_{T}$ for the target domain $T$. $g_{\psi}$ (e.g., the Dino linear head or the MAE decoder) is discarded since it is not part of the ViT encoder $f$. Post-ExPLoRA, only $\Delta_{T}$, consisting of 1-2 unfrozen ViT blocks, LoRA matrices, and layer-normalization weights, are stored for each $T$– all of which can be merged into the original ViT, thus preserving architecture. Like LoRA, ExPLoRA significantly reduces additional storage requirements compared to fully training $W_{T}$ from scratch.
We can treat $W_{T}^{\ast}$ as any other fully pre-trained ViT $W_{T}$. Finally, solving for $\Delta^{(\tau)}$ on each labeled dataset ${\mathcal{Y}}^{(\tau)}_{T}$:
$\Delta^{(\tau)}=\operatorname*{arg\,min}_{\begin{subarray}{c}\phi\\ \theta\in\Theta(r)\end{subarray}}\sum_{({\mathbf{x}},{\mathbf{y}})\in{\mathcal{Y}}^{(\tau)}_{T}}{\mathcal{C}}^{\prime}\left(h^{(\tau)}_{\phi}\left(f_{\theta}({\mathbf{x}};W_{T}^{\ast})\right),{\mathbf{y}}\right)$ (6)
where the ViT $f_{\theta}$ is now initialized with $W_{T}^{\ast}$, $h^{\tau}_{\phi}(\cdot)$ is a task-specific ViT decoding head with parameters $\phi$, and ${\mathcal{C}}^{\prime}$ represents a supervised-learning cost function.
For any PEFT method, we restrict $\theta\in\Theta(r)$ where $r$ controls the fraction of trainable weights in ${\mathcal{L}}$. (e.g., LoRA only on attention $Q,V$ matrices). For linear probing, $\Theta(0)=\emptyset$, so $(h^{\tau}_{\phi}\circ f_{\theta})({\mathbf{x}};W_{T}^{\ast})=\phi^{\top}f({\mathbf{x}};W_{T}^{\ast})$. For general fine-tuning, we optimize both $\theta\in\Theta(\infty)$ and $\phi$, modifying $h^{\tau}$ as per the task (e.g., a decoder head for segmentation). With $\Delta^{(\tau)}$, eq. 4 gives us our final model weights $W^{(\tau)}_{T}$, which can be used for classification, segmentation, detection etc.
@@ 6 Experiments
Our experimental results consist of a case study on satellite imagery (section 6.1), with an ablation study in section 6.1.2. We evaluate on multiple downstream tasks in sections 6.1.3, 6.1.4 and 6.2. Additional experiments and ablations are provided in appendix B and training hyperparameter and compute configurations are mentioned in appendix C. Our results achieve a new SoTA top 1 accuracy of 79.3% ($\uparrow$1.5%) on the competitive fMoW-RGB benchmark, outperforming fully pre-trained and fine-tuned models while using 6% of the ViT encoder parameters. We also achieve a $\uparrow$8.2improvement in linear probing accuracy on the same dataset. Across other satellite datasets, we match fully-pretrained prior state-of-the-art methods that required up to 8x more compute and 16x more trainable parameters and demonstrate competitive performance on WiLDS benchmark datasets as well.
@@ 6.1 Case Study: Satellite Imagery
We examine satellite images given their importance towards societal applications (Broader Impact) and since they represent a significant domain shift from natural images. There is a large and growing body of research on developing foundation models for satellite imagery from scratch (Cong et al., 2022; Reed et al., 2023; Tang et al., 2024), thus presenting a good benchmark for ExPLoRA.
@@ 6.1.1 RGB Satellite Images
We first consider the functional map of the world (fMoW) dataset of high-resolution satellite images, each paired with one of 62 classification labels (Christie et al., 2018). fMoW is used as a benchmark for satellite-image foundation models (Cong et al., 2022; Reed et al., 2023).
Model PEFT
Pre-train/Fine-tune
@@
ScaleMAE [54] Full 303.3M / 303.3M 77.80
SatMAE [17] Full 303.3M / 303.3M 77.78
SatMAE [17] LoRA-$r8$ [28] 303.3M / 0.8M 76.10
ScaleMAE [54] LoRA-$r8$ [28] 303.3M / 0.8M 78.01
GFM [44] LoRA-$r8$ [28] 303.3M / 0.8M 73.03
GDA [56] GDA-$r16$ [56] 8.5M / 8.5M 71.88
MAE [26] LoRA-$r8$ [28] - / 0.8M 76.21
M-[L]-$r64$ LoRA-$r8$ [28] 18.7M / 0.8M 76.55
DinoV2 [47] LoRA-$r8$ [28] - / 0.8M 78.08
DinoV2 [47] BOFT-$b2m8$ [40] - / 0.9M 72.40
DinoV2 [47] Mona [69] - / 7.1M 72.80
DinoV2 [47] VPT-100 [29] - / 0.4M 77.29
DinoV2 [47] GVPT-100 [70] - / 0.4M 76.22
DinoV2 [47] AdaLoRA-$r8$ [72] - / 1.2M 78.87
DinoV2 [47] Adapter+ [59] - / 1.4M 78.16
DinoV2 [47] SA2VP [49] - / 1.1M 77.53
D-[L]-$r64$ SA2VP [49] 18.7M / 1.1M 78.51
D-[L]-$r64$ LoRA-$r8$ [28] 18.7M / 0.8M 79.28
Table 1: Results on the fMoW-RGB validation dataset. The “Pre-train / Fine-tune #Params" refer to the trainable parameters of the ViT-L encoder required on the new domain, i.e. satellite images. M-[L]-$r64$ and D-[L]-$r64$ refer to ExPLoRA models initialized with MAE and DinoV2 weights, respectively (section 5).
Method Arch.
Top 1 Acc.
@@
Table 2: Linear-probing on fMoW-RGB. The first four rows fully pre-train on the dataset. $\dagger$ denotes concatenating features from the last 4 ViT blocks. All other rows use features of the last ViT block.
We compare our results in table 1 against both prior fully pre-trained SoTA foundation models as well as PEFT techniques applied on ViTs pre-trained with MAE and/or DinoV2 weights. Our results demonstrate that D-ExPLoRA-$\tt{[L]}$-$r64$ is SoTA in terms of fMoW-RGB average accuracy at 79.28%. ExPLoRA outperforms techniques that require fully and/or continually pre-training ViTs on fMoW while using 6% of the original ViT encoder parameters. Further experiments with MAE are in B.6.
ExPLoRA-initializations with LoRA fine-tuning outperform other unsupervised initializations paired with PEFT techniques by 1-3%, including SoTA matrix-adaptation methods like AdaLoRA (Zhang et al., 2023b), BOFT (Liu et al., 2023), VPT approaches such as GVPT (Yoo et al., 2023), SA2VP (Pei et al., 2024), and adapter methods like Adapter+ (Steitz & Roth, 2024). We also outperform satellite image continual pre-training methods such as GFM (Mendieta et al., 2023) and GDA (Scheibenreif et al., 2024) by 6%. Additionally, applying SA2VP to ExPLoRA-initialized ViTs improves performance over DinoV2 by  1%, showcasing ExPLoRA’s compatibility with other PEFT methods and its versatility as an initialization for new domains.
6.1.2 Ablation study
@@
We perform an ablation study (table 3) on linear-probing performance for fMoW-RGB to determine whether our proposed configuration performs optimally.
@@
Next, we investigate whether high LoRA ranks used on all ViT layers (i.e. all attention and MLP matrices, not just $Q,V$) is beneficial. Surprisingly, this significantly harms learning (row 4, 5). In fact, it is much less effective than using just LoRA-$r256$ on the $Q,V$ matrices of all ${\mathcal{L}}$ blocks (row 3). However, both rows 3 and 4 are much less parameter-efficient than ExPLoRA (rows 6-8, 11-13).
@@
Lastly, we investigate the impact of LoRA rank on ExPLoRA. Changing the rank from 8 to 32 has a small improvement ($\uparrow 0.09\%$), but changing from 32 to 64 brings about a much larger improvement ($\uparrow 1.08\%$), with only a relatively small increase in trainable parameters. This demonstrates that higher ranks are necessary during pre-training for effective learning on the new domain. Further ablations on compute efficiency (B.2), data efficiency (B.3), MAE decoder rank (B.4), and ViT backbone size (B.5) are in appendix B.
@@ 6.1.3 Multi-Spectral Satellite Images
Model PEFT
Pre-train/Fine-tune
@@
Top 1
Acc.
MAE [26] Full - / 303.3M 51.61
SatMAE [17] Full 303.3M / 303.3M 61.48
MAE [26] LoRA-$r8$ / 0.8M 46.97
SatMAE [17] LoRA-$r8$ 303.3M / 0.8M 59.48
GFM [44] LoRA-$r8$ 303.3M / 0.8M 57.55
GDA [56] GDA-$r16$ 7.3M / 7.3M 55.23
[…]
M-[L]-$r32$ LoRA-$r8$ 16.2M / 0.8M 51.84
M-[1,L]-$r32$ LoRA-$r8$ 29.7M / 0.8M 60.15
[…]
@@
Table 5: Classification results on the validation set of fMoW-Temporal. All MAE/SatMAE experiments use ViT-L, while GASSL uses a ResNet.
@@
Table 6: Results on the validation sets of SpaceNet-v1 (SN), a segmentation task, and Resisc-45 (R45), a classification task.
@@
Table 9: Object detection results on the validation set of GlobalWheat. AP and AR stand for average precision and average recall.
Dataset
Next, we consider the fMoW-Sentinel dataset, a large dataset of Sentinel-2 images introduced by Cong et al. (2022). Each image consists of 13 spectral bands and is paired with one of 62 classes.
With fMoW-Sentinel, we evaluate transfer from natural images to multi-spectral, low-resolution satellite images - a harder task than fMoW-RGB due to the absence of non-RGB bands in $S$. We use the group-channel ViT-L from Cong et al. (2022), initialized with MAE. During algorithm 1, we additionally unfreeze only the patch embedding layers due to architectural differences.
Table 6 shows the challenge: fully fine-tuning from MAE drops accuracy by nearly 10% (row 1 vs row 2), LoRA tuning from MAE performs worse (row 4), and unfreezing four transformer blocks (row 6) fails to help. However, ExPLoRA with ${\mathcal{U}}=\{\tt{1,L}\}$ outperforms even full pre-training from scratch for LoRA fine-tuning (row 5 vs. last row), demonstrating effective adaptation to a very different domain while using <10% of the parameters.
6.1.4 Additional Satellite Datasets
We perform extensive experiments on downstream satellite datasets, with further results in B.1.
fMoW-Temporal
Each input is a sequence of up to 3 fMoW-RGB (Christie et al., 2018) images of a location, distributed temporally, and paired with one of 62 classes. Since the inputs are now temporal sequences, we initialize the temporal MAE architecture from Cong et al. (2022) with MAE weights, and pre-train on ${\mathcal{X}}_{T}$ with ${\mathcal{U}}=\tt[L]$ and LoRA rank 32. ExPLoRA then outperforms temporal SatMAE for PEFT (table 6), demonstrating successful transfer learning at a fraction of the pre-training parameters.
SpaceNet-v1
RESISC-45
The RESISC-45 (Cheng et al., 2017) benchmark dataset consists of 31,500 satellite images of varying resolution (0.2m-30m GSD), with 45 classes. The data is split into 25,200 training and 6,300 validation images, as per Reed et al. (2023). In table 6, our D-ExPLoRA pre-trained on only high-resolution fMoW-RGB images (last row of table 1) achieves SoTA results of 97.32% on multi-resolution RESISC-45 images, with just linear-probing. We demonstrate successful transfer learning from ExPLoRA pre-training, without requiring any additional modifications for scale-aware representation learning (Reed et al., 2023).
6.2 WiLDS Datasets
We test ExPLoRA on the WILDS (Koh et al., 2021) benchmark, specifically on Camelyon17 (Bandi et al., 2018), iWildcam (Beery et al., 2020) and GlobalWheat (David et al., 2020, 2021) datasets, representing domain transfers to medical, wildlife, and agricultural imagery, respectively.